OpenStudy (anonymous):

Find an equation of the line tangent to the graph of y = sin(x) + 3 at x = π.

OpenStudy (jhannybean):

1. find \(\ f'(x)\) 2. plug in \(\ x=\pi\) into \(\ f'(x)\). This will give you your slope 3. Use \(\ x = \pi\) to find y-value. \(\ f(\pi)= y_0\). This will give you your coordinate (\(\pi , y_0)\) 4. Input all these values into the point-slope formula \(\ y - y_0 = f'(x)(x-x_0)\) 5. write formula in slope-intercept form : \(\ y=f'(x)( x) +y_0\)

OpenStudy (anonymous):

thank you